In a trial of 123 patients who received​ 10-mg doses of a drug​ daily, 32 reported headache as a side effect. Obtain a point estimate for the population proportion of patients who received​ 10-mg doses of a drug daily and reported headache as a side effect.

P^=
Answer #1

Solution :

Given that,

Point estimate = sample proportion = = x / n = 32 / 123 = 0.260

= 0.260